# Privacy in Speech Technology

### v3

@@ Abstract
Speech technology for communication, accessing information, and services has rapidly improved in quality. It is convenient and appealing because speech is the primary mode of communication for humans. Such technology, however, also presents proven threats to privacy. Speech is a tool for communication and it will thus inherently contain private information. Importantly, it however also contains a wealth of side information, such as information related to health, emotions, affiliations, and relationships, all of which are private. Exposing such private information can lead to serious threats such as price gouging, harassment, extortion, and stalking. This paper is a tutorial on privacy issues related to speech technology, modeling their threats, approaches for protecting users’ privacy, measuring the performance of privacy-protecting methods, perception of privacy as well as societal and legal consequences. In addition to a tutorial overview, it also presents lines for further development where improvements are most urgently needed.
@@ I Introduction
Speech is a mode of communication and thus inherently contains a wealth of information (see table I). In addition to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information. Moreover, speaking is a dynamic interaction between two or more speakers. Dialogues thus also contain information about the relationship between participants, such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history. From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [1, 2, 3, 4].
@@
The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression damaging democratic societies. Users may feel they lose control of personal aspects of their lives.
@@
Studying and improving privacy in speech technology is important because breaches in privacy can have serious consequences. Table II lists examples of threats and exploits. The list is incomplete, and we can expect new threats to be discovered. The generic solution template is, however, typically always the same: Minimize transmission and storage of as well as access to sensitive information that is irrelevant to the service that the users want. It is then ”merely” a question of how such minimization is achieved, what information is relevant, and how to determine what the user wants (provide information and enable control of services and threats).
@@
With respect to threats, media attention is often focused on breaches which have large economic consequences [5, 6, 7]. While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways. First, a breach is a worst-case event, whereas threats that have not yet led to a breach event can already have a large impact. For example, users may choose not to use systems that threaten their privacy; known weaknesses and even a lax attitude of the service provider toward privacy can, therefore, have an effect on the adoption, retention, and sales of products and services. Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat. To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
@@
Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [8]. With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II). While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large and their prevalence makes them a considerable threat also economically.
@@
This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms. The range of applications and content types where privacy-preserving processing methods are needed is vast, and section II presents a brief overview thereof, and section III defines the vocabulary. Analysis of threats demonstrates that while attack surfaces have great variety, the threat models are similar (see section IV). Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V). To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section VI). While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level. For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VII). Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
@@
This paper focuses primarily on speech technology in human-computer interfaces (see fig. 1), but will also address telecommunication where appropriate. Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g. [9]). Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [10, 11, 12].
@@
To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology. Recent works in the field, however, include a technical overview [13], a survey on privacy issues with voice assistants [14], as well as a popular-science review [15] of this area, and several doctoral theses have their respective summaries, e.g. [16, 17, 18, 19, 20, 21, 22, 23]. Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [9], in statistical theory of privacy [24], in social sciences [25] and psychology [26]. There is even an excellent and thorough meta-study of all areas of science that discuss privacy [27]. Central contributions of this work include 1. structuring and collecting information, research results, and vocabulary into a unified format, 2. proposing a framework for analysis of attack vectors (see fig. 1), and 3. an introduction to the psychological aspects of privacy that is not often discussed in engineering-oriented publications (see section VII).
@@ II Content categories and applications
To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion both on categories of private information, complementing table I as well as the application of such information. Observe that this is not a complete list (nor is table I) of the content nor of application categories. The purpose is merely to provide a characterization of work done and challenges in the research field, with accompanying references.
@@
Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy. We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g. [28, 37, 44, 44, 45, 46, 47, 48, 49, 29, 30, 31, 32, 46, 47, 33, 34, 35, 36, 38, 39, 40, 41, 42, 43]. Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g. [50, 51, 52, 53, 54, 55, 44, 56, 57]. Speaker characterization is the natural complement of speaker idenfication [13, 58]. Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming. Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
@@
Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research. With respect to privacy, it contains two obvious challenges for privacy. We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g. [59, 60, 61, 62, 63, 64, 65], or we can use natural language processing to anonymize the text content, e.g. [66, 67, 68, 69, 9, 70].
@@
Privacy is even more important in always-on applications like wake-word detection, i.e. when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”. This application is more sensitive exactly because it is always on as users cannot choose when data is processed [71, 72, 73, 74]. The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [75, 76]. Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [72, 77, 78, 30, 79, 80, 81].
@@
Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [82]. This task can have an impact on privacy in two primary ways. First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [83, 84]. Attenuating or removing background noises and competing speakers, as well as removing reverberation, can thus improve privacy. In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [80, 18, 85].
@@
Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [86, 87, 13, 88, 89]. Some, however, limit the term biometric to refer only to physical and behavioral characteristics. Such properties include health [90], emotions [91] and gender identity [92, 93, 73, 46], each warranting their own treatment.
@@
Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows). Each threat is marked with the number of the corresponding part of section IV-D. Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
@@ III-A Privacy and security
Privacy is closely related to challenges in security, and it is often difficult to distinguish between them. Here, information security refers to enforcing rightful access to information in scenarios where the access rights are binary (granted or not granted). In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [94]). The rules may distinguish between the ability to access and de facto accessing the data, different ways and purposes of using the data, and sharing the data with third parties. The rules may further be explicit (e.g., verbalized or documented) or implicit (as based on, e.g., social norms and tacit expectations) and can include requirements on monitoring, reporting, and control. For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig. 1). Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
@@ III-B Information categories
Target information is the content of a communication that is required to fulfill the purpose of said communication. Side-information (also known as auxiliary information, external knowledge, or background knowledge [95]) is any and all information in a communication other than the target information. See fig. 2.
Attributes (or characteristics [4]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features. Table I lists examples of speech attributes with private information.
Observe that while the target information is often the lexical information contained in speech, in many applications, the target information can be some other attribute. For example, in medical analysis, the target information is the health state of the speaker. In that case, the lexical information may be private side-information.
@@
Figure 2: The high-level threat model of speech interaction, where a target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message. It is a threat to privacy when an undesired recipient gains access to that target information or side information (marked by red arrows and exclamation marks).
@@ III-C Tasks and processing objectives
When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer, and this is known as the trusted task. In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform other nefarious tasks. [47, 92].
@@
Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [96]. Re-identification refers to identifying individuals from apparently anonymized information [88, 97].
In anonymization tasks, enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s). It is used in the context of trusted tasks, where the agent is authorized to access this information. In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data. Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data. [98]

### v4

@@ Abstract
Speech technology for communication, accessing information, and services has rapidly improved in quality. It is convenient and appealing because speech is the primary mode of communication for humans. Such technology, however, also presents proven threats to privacy. Speech is a tool for communication, and it will thus inherently contain private information. Importantly, it also contains a wealth of side information, including details about health, emotions, affiliations, and relationships, all of which are private. Exposing such private information can lead to serious threats such as price gouging, harassment, extortion, and stalking. This paper is a tutorial on privacy issues related to speech technology, modeling their threats, approaches for protecting users’ privacy, measuring the performance of privacy-protecting methods, perception of privacy, as well as societal and legal consequences. In addition to a tutorial overview, it also presents lines for further development where improvements are most urgently needed.
@@ I Introduction
Speech is a mode of communication and thus inherently contains a wealth of information (see table I). In addition, to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information. Moreover, speaking is a dynamic interaction between two or more speakers. Dialogues thus also contain information about the relationship between participants, such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history. From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [singh2019profiling, kroger2019privacy, kroger2022personal, singh2025humanvoiceunique].
@@
The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression, damaging democratic societies. Users may feel they lose control of personal aspects of their lives.
@@
Studying and improving privacy in speech technology is important because breaches in privacy can have serious consequences. Table II lists examples of threats and exploits. The list is incomplete, and we can expect new threats to be discovered. The generic solution template is, however, typically always the same: Minimize transmission, storage, and access to sensitive information that is irrelevant to the service that the users want. It is then ”merely” a question of how such minimization is achieved, what information is relevant, and how to determine what the user wants (provide information and enable control of services and threats).
@@
With respect to threats, media attention is often focused on breaches which have large economic consequences [Lynskey2019GuardianPrivacyScandal, hern2019, brewster2021]. While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways. First, a breach is a worst-case event, whereas threats that have not yet led to a breach event can already have a large impact. For example, users may choose not to use systems that threaten their privacy; known weaknesses and even a lax attitude of the service provider toward privacy can, therefore, have an effect on the adoption, retention, and sales of products and services. Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat. To maintain users’ trust, it is therefore essential to both uphold the users’ actual privacy but also design systems that actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
@@
Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [levy2020intimate]. With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II). While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large, and their prevalence makes them a considerable threat also economically.
@@
This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms. The range of applications and content types where privacy-preserving processing methods are needed is vast, and section II presents a brief overview thereof, and section III defines the vocabulary. Analysis of threats demonstrates that while attack surfaces have a great variety, the threat models are similar (see section IV). Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private information through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V). To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section VI). While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level. For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VII). Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which have motivated governments to increasingly regulate the use of technology (see section VIII).
@@
This paper focuses primarily on speech technology in human-computer interfaces (see fig. 1), but will also address telecommunication where appropriate. Focus is here, however, limited to the acoustic speech signal, also known as the voice, since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g. [mahendran2021privacy]). Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended also to cover mobile telecommunication and voice-over-Internet protocols (VoIP) [penders2004privacy, wong2009privacy, arapinis2017analysis].
@@
To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology. Recent works in the field, however, include a technical overview [nautsch2019preserving], a survey on privacy issues with voice assistants [acosta2022survey], as well as a popular-science review [backstrom2020privacy] of this area, and several doctoral theses have their respective summaries, e.g. [Nelus2022, das2021thesis, zarazaga2022thesis, pathak2012privacy, williams2022learning, portelo2015privacy, vestman2020phd, nespoli2025phd]. Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [mahendran2021privacy], in statistical theory of privacy [cummings2023challenges], in social sciences [petronio2002boundaries] and psychology [Derlega1977PrivacyAS]. There is even an excellent and thorough meta-study of all areas of science that discuss privacy [van2018handbook]. Central contributions of this work include 1. structuring and collecting information, research results, and vocabulary into a unified format, 2. proposing a framework for analysis of attack vectors (see fig. 1), and 3. an introduction to the psychological aspects of privacy that is not often discussed in engineering-oriented publications (see section VII).
@@ II Content categories and applications
To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion on both categories of private information, complementing table I as well as the application of such information. Observe that this is not a complete list (nor is table I) of the content nor of application categories. The purpose is merely to provide a characterization of work done and challenges in the research field, with accompanying references.
@@
Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy. We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g. [chouchane21_interspeech, granqvist20_interspeech, mawalim20_interspeech, mawalim20_interspeech, nautsch19b_interspeech, nelus19_interspeech, nelus19b_interspeech, noe21_interspeech, perero2022x, zewoudie2021federated, zewoudie2021use, teixeira22_interspeech, treiber2019privacy, nelus19_interspeech, nelus19b_interspeech, BAI202165, han2020voice, KINNUNEN201012, hansen2015speaker, parthasarathi2013wordless, pathak11_interspeech, parthasarathi11_interspeech, pathak2012ppsa, jimenez2017two, portelo2014bprivacy]. Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g. [noe20_interspeech, pierre22_interspeech, prajapati21_interspeech, shamsabadi2023diff, srivastava20_interspeech, Srivastava2022taslp, mawalim20_interspeech, maouche20_interspeech, Fang2019]. Speaker characterization is the natural complement of speaker idenfication [nautsch2019preserving, nguyen2010automatic]. Such methods related to speaker identity can be applied, for example, to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming. Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
@@
Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research. With respect to privacy, it presents two obvious challenges. We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g. [shoemate2022sottovoce, nguyen2023federated, aloufi21_interspeech, garg20b_interspeech, shah21_interspeech, srivastava19_interspeech, tomashenko2022privacy], or we can use natural language processing to anonymize the text content, e.g. [liu22_interspeech, maouche2022enhancing, adelani20_interspeech, novotney21_interspeech, mahendran2021privacy, habernal2021differential].
@@
Privacy is even more critical in always-on applications like wake-word detection, i.e., when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”. This application is more sensitive exactly because it is always on, as users cannot choose when data is processed [koppelmann21_interspeech, koppelmann22_interspeech, stoidis22_interspeech, yang21_interspeech]. The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [haider19_interspeech, salah2025privacy]. Such ambient voice interfaces are often implemented through acoustic sensor networks, which pose their own challenges [koppelmann22_interspeech, oconnor20_interspeech, malekzadeh2018replacement, zewoudie2021use, zarazaga2020acoustic, nelus2021, backstrom2025privacy].
@@
Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [benesty2008springer]. This task can have an impact on privacy in two primary ways. First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [zarazaga21_interspeech, rech2022multi]. Attenuating or removing background noises and competing speakers, as well as removing reverberation, can thus improve privacy. In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [nelus2021, zarazaga2022thesis, hendriks2013ppdistributed].
@@
Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [adler2006towards, nautsch20_interspeech, nautsch2019preserving, ISObiometric, edpb2021guidelines]. Some, however, limit the term biometric to refer only to physical and behavioral characteristics. Such properties include health [natarajan2021prioris], emotions [feng22b_interspeech] and gender identity [nelus2018itg, stoidis21_interspeech, stoidis22_interspeech, nelus19_interspeech], each warranting their own treatment.
@@
Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network, or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows). Each threat is marked with the number of the corresponding part of section IV-D. Dotted red arrows marked with “X” indicate threats outside the current scope, and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
@@ III-A Privacy and security
Privacy is closely related to challenges in security, and it is often difficult to distinguish between them. Here, information security refers to enforcing rightful access to information in scenarios where the access rights are binary (granted or not granted). In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [nissenbaum2004privacy]). The rules may distinguish between the ability to access and de facto accessing the data, different ways and purposes of using the data, and sharing the data with third parties. The rules may be explicit (e.g., verbalized or documented) or implicit (based on social norms and tacit expectations), and can include requirements for monitoring, reporting, and control. For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig. 1). Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
@@ III-B Information categories
Target information is the content of a message that is required to fulfill the purpose of that communication. Side-information (also known as auxiliary information, external knowledge, or background knowledge [ganta2008composition]) is any and all information in a message other than the target information. See fig. 2.
Attributes (or characteristics [singh2025humanvoiceunique]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style, or environment, the recording configuration, or any other related features. Table I lists examples of speech attributes with private information.
Observe that while the target information is often the lexical information contained in speech, in many applications, the target information can be some other attribute. For example, in medical analysis, the target information is the health state of the speaker. In that case, the lexical information may be private side information.
@@
Figure 2: The high-level threat model of speech interaction, where target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message. It is a threat to privacy when an undesired recipient gains access to that target information or side information (marked by red arrows and exclamation marks).
@@ III-C Tasks and processing objectives
When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer, and this is known as the trusted task. In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform other nefarious tasks. [nelus19b_interspeech, nelus2018itg].
@@
Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [gong2018attribute]. Re-identification refers to identifying individuals from apparently anonymized information [ISObiometric, ISOIEC224745_2022].
In anonymization tasks, enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s). It is used in the context of trusted tasks, where the agent is authorized to access this information. In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data. Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data. [tomashenko2022voiceprivacy]